In a survey of a Tridelt Chapter with 50 members, 19 were taking mathematics, 18 were taking English, and 9 were taking both. How many were not taking either of these subjects?
1
Expert's answer
2013-01-17T09:33:29-0500
Solution (1):
19-9=10 members were taking only mathematics
18-9=9 members were taking only English
10+9+9=28 members were taking only mathematics and only English and only both
50-28=22 members were not taking either of these subjects
Answer: 22 members
Solution (2):
19+18=37 members were taking mathematics and English
37-9=28 members& were taking only mathematics and only English
50-28=22 members were not taking either of these subjects
Answer: 22 members.